OpenStudy (anonymous):

How can I find the vertex, focus, directrix and focal width of a parabola with the equation -1/20x^2=y?

OpenStudy (jack1):

do you know what each of those things are...?

OpenStudy (anonymous):

Yes! How do I determine them from the equation though?

OpenStudy (jack1):

so your vertex is the turning point of your graph, yeah?

OpenStudy (anonymous):

Yes.. It would be at the origin in this case

OpenStudy (jack1):

so how did you want to work it out... you could either draw the graph and get your answers from there, or you can do it mathematically

OpenStudy (anonymous):

Id like to do it mathematically, so that I can do it without a calculator in case I ever have to. If you could help me figure "P" out then I think I can do it by myself.

OpenStudy (jack1):

and your equation is...? -(1/20) x^2=y or -1/(20x^2) =y ???

OpenStudy (anonymous):

The first

OpenStudy (jack1):

k, well that can be re-written as -x^2 / 20 = y so the turning poing of a parabola occurs when the gradient = 0 so what you can do is take the derivative (which give the gradient at any x point) and sub in y=0

OpenStudy (jack1):

so derivative is y = -x/10 so sub in 0 so 0 = -x/10 so x = 0 turning point is a 0,0

OpenStudy (anonymous):

So when substitute for that, the answer will be P? In my class our formula for a vertical parabola is x^2=4py

OpenStudy (jack1):

...never heard of P, sorry

OpenStudy (jack1):

what does it let u work out?

OpenStudy (anonymous):

Basically everything. For example, if p>0 and the parabola opens downward, then the focus is (h,k+p) and the directrix is y=k-p

OpenStudy (anonymous):

Oh! Okay, thank you

OpenStudy (jack1):

so your focus is (0, -4) ...i think

OpenStudy (phi):

you start with \[ -\frac{1}{20} x^2=y \] multiply both sides by -20 : \[ -20 \cdot -\frac{1}{20} x^2= -20y \\ x^2 = -20y\] factor out a 4 from the 20 (i.e. write 20 as 4 * 5): \[ x^2 = 4(-5)y \] as you can see p=-5

OpenStudy (anonymous):

Thank you guys!

OpenStudy (jack1):

for the focus, switch the terms so you have x = (something something) y in this case x^2 = -20y so x^2 = 4 *(-5)y P = -5 (sorry, typo before)

OpenStudy (anonymous):

No problem!

OpenStudy (jack1):

oooh, shiny!